OpenStudy (anonymous):

Find the value of 10 times the square root of 2 over the square root of 5..

OpenStudy (hba):

\[10*\sqrt{2}/\sqrt{5}\]

OpenStudy (saifoo.khan):

\[\frac{10 \times \sqrt2}{\sqrt5} \times \frac{\sqrt5}{\sqrt5}\]

OpenStudy (anonymous):

rationalize the denominator by writing \[\frac{10\sqrt{2}}{\sqrt{5}}\times \frac{\sqrt{5}}{\sqrt{5}}\] oh hello @saifoo.khan

OpenStudy (saifoo.khan):

Hello @satellite73

OpenStudy (anonymous):

saifoooooooooooooooooooooooooooooooo

OpenStudy (anonymous):

@Ashley16 is it clear how to finish this?

OpenStudy (saifoo.khan):

http://t.qkme.me/3q88p4.jpg

OpenStudy (saifoo.khan):

http://t.qkme.me/7cd5.jpg

OpenStudy (anonymous):

lolololol!!

OpenStudy (anonymous):

yess, Thank you! (:

OpenStudy (anonymous):

yw, don't forget to cancel

OpenStudy (anonymous):

got it!